OpenStudy (senia01):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –4 3 –2 4 Line 2 x y 0 8 2 6 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (senia01):

@q12157

OpenStudy (anonymous):

(4-3)/(-2-(-4))=1/2

OpenStudy (anonymous):

6-8/2-0=-1

OpenStudy (anonymous):

Different slopes=they must intersect=exactly 1

OpenStudy (senia01):

Thank you so much

OpenStudy (anonymous):

:3